# Neutrinos from explosive transients at the dawn of multi-messenger astronomy

### v1

@@ 2.1.3 Detection strategies for multi-messenger searches
Among existing neutrino telescopes, the ones allowing for the largest supernova neutrino event rate are Super-Kamiokande 84 and the IceCube Neutrino Observatory 85 (cf. Table 1 of 86 for an overview on neutrino telescopes and expected event statistics). While for core-collapse supernovae in our Galaxy, IceCube would guarantee the largest statistics, Super-Kamiokande would be background free with good approximation, given the small time window characterizing the supernova neutrino burst and should allow us to reconstruct event-by-event energy information.
Within the multi-messenger framework, the SuperNova Early Warning System (SNEWS) is an operating global network of neutrino telescopes that aims to release a prompt alert of core-collapse events to enable multi-messenger observations 87. In addition to being part of SNEWS, both Super-Kamiokande and IceCube are heavily involved in the development of multi-messenger alert systems. For example, Super-Kamiokande, in collaboration with the KamLAND detector, has established an alert system that searches for pre-supernova neutrinos and could release a warning about an upcoming electromagnetic burst about a dozen hours earlier 88. The IceCube supernova data acquisition system is being updated to swiftly react to alerts from the LIGO-Virgo-KAGRA gravitational wave network 89.
Hyper-Kamiokande 90 is expected to begin taking data in 2027 and the liquid scintillator Jiangmen Underground Neutrino Observatory (JUNO) 91 should be operative already in 2025. These neutrino telescopes will guarantee large event statistics of supernova neutrinos with competitive pointing capabilities. If able to detect supernova neutrinos, the liquid-Argon Deep Underground Neutrino Experiment (DUNE) 92 could provide key complementarity with respect to Hyper-Kamiokande and JUNO, being mostly sensitive to electron neutrinos rather than electron antineutrinos. In addition, experiments for the direct detection of dark matter, exploiting the coherent elastic scattering of neutrinos on nuclei, such as XENONnT 93 at the Gran Sasso National Laboratory and the planned neutrino detector RES-NOVA 94, have the potential to complement the potential of large-scale neutrino detectors, being sensitive to all neutrino flavors.
@@ 2.1.4 Supernova physics signatures imprinted in the neutrino signal
Despite the dependence of the neutrino signal on the progenitor mass of the collapsing star as well as the nuclear equation of state, neutrino detection will be instrumental in exploring the following 32; 78.
  • Alert of the stellar collapse. Neutrinos reach Earth up to one day earlier than electromagnetic radiation. Pre-supernova neutrinos 95 (cf. 96 for a summary of the thermal processes leading to neutrino production during the stages of stellar evolution of a massive star) and the early detection of the supernova neutrino burst could inform astronomers of the upcoming electromagnetic burst, enabling the electromagnetic detection of the shock breakout 97. We can also use the neutrino signal to localize the stellar collapse in the sky and define the time window to search for gravitational waves as well as build a matching filter for gravitational wave searches 98; 99.
  • Black hole forming collapses. The temporal evolution of the neutrino signal carries distinctive signatures, ending abruptly at the time of black hole formation 100. Such collapses would not be detectable through electromagnetic radiation. Therefore, the detection of neutrinos and gravitational waves could be essential to assess the rate of black hole forming collapses, complementing ongoing electromagnetic searches 101.
  • Pre-explosion physics. The temporal evolution of the neutrino event rate (during the accretion phase) carries clear imprints of the hydrodynamical instabilities characterizing the accretion phase, such as SASI and convection 102. Rotation can also affect the neutrino signal 103; 104.
  • Neutron star properties. The late-time neutrino emission (during the Kelvin-Helmoltz cooling phase) strongly depends on the nuclear equation of state and the neutron-star properties. Despite the presence of degeneracies in the neutrino signal, its detection could place constraints on the physics of the compact remnant 105; 106.
  • Physics beyond the Standard Model. The presence of exotic signatures in the time evolution of the neutrino signal and/or its energy distribution could be a manifestation of New Physics 107. However, degeneracies among multiple New Physics scenarios exist and are yet to be fully understood 108. According to the type of exotic features eventually appearing in the observed neutrino signal, before to draw any conclusions on the discovery of New Physics relying on core-collapse supernovae, one would need to model the standard core-collapse supernova scenario (including flavor conversion) consistently–an ambitious goal yet to be achieved.
@@ 2.1.5 Diffuse supernova neutrino background
Every second a supernova explodes somewhere in the Universe. The cumulative flux of neutrinos from all core-collapse supernovae is named diffuse supernova neutrino background (DSNB) 109. This stationary and isotropic flux depends on the core-collapse supernova rate as a function of redshift, $R(z)$, and the time-integrated (anti)neutrino spectral energy distribution $F_{\nu}(E)$ emitted from each collapsing star:
@@
To date, the largest uncertainty plaguing the DSNB is the core-collapse supernova rate 110, while the DSNB negligibly depends on the uncertainties on the initial mass function 111. The unknown rate of collapses leading to black hole formation 112 and magnetorotational collapses 113, as well as the metallicity evolution of the host galaxies 114, and the impact of mass transfer during binary evolution (if the collapsing star happens to be in a binary system) 115 are some of the uncertainties in the forecast of the DSNB. In addition, the unknowns linked to the supernova microphysics (such as the neutron star mass, high-density nuclear equation of state, and the flavor conversion physics 116; 117) also affect the DSNB.
The upcoming DSNB detection will be instrumental in placing neutrino-based constraints on the properties of the supernova population, the average redshift distribution of supernovae in the local Universe, and the average supernova energetics. In particular, the rate of core-collapse supernovae measured through electromagnetic surveys could be biased because electromagnetically dark collapses could be missed, but such collapses would contribute to the DNSB 118.
The Gadolinium loading of the water Cherenkov detector Super-Kamiokande has drastically reduced the signal over background ratio, already excluding some of the most optimistic DSNB models as well as hinting towards a signal excess at $2.3\sigma$ 119; 64. The latest Super-Kamiokande upper limits on $\bar{\nu}_{e}$’s of astrophysical origin are presented in the right panel of Fig. 1 together with the expected DSNB signal. In the future, Hyper-Kamiokande 90, possibly loaded with Gadolinium, and JUNO 120 will allow us to investigate the DSNB physics with larger statistics (cf. Table 2 of 113 for estimations of the expected neutrino number of events). On the other hand, neutrino and direct-detection dark matter experiments, exploiting the coherent elastic neutrino scattering on nuclei, can place upper limits on the DSNB emission in the non-electron flavors 121.
@@ 2.2.1 Neutron-star merger physics
In 2017, the joint detection of gravitational waves and electromagnetic radiation from the merger of two neutron stars confirmed the relevance of these objects for astrophysics and fundamental physics 4; 122; see Table 1 for a summary of the multi-messenger energetics. The detection in 2017 has provided a strong indication that at least a fraction of the elements heavier than iron can be synthesized in these sources. Yet, several aspects of the physics of merger remnants remain to be clarified, such as the interplay between magnetic fields and neutrino transport, and their degree of sophistication 123; 124.
@@
Following the merger event, the accretion disk undergoes cooling via neutrino emission, with neutrino-matter interactions determining the composition of the merger outflows 125. According to the mass of the system and the nuclear equation of state, the central compact object (a newly-born neutron star) can immediately collapse into a black hole or it can be a rapidly spinning neutron star remnant. In the latter scenario, a hyper-massive 126 or supra-massive 127 neutron star may form, which is expected to collapse into a black hole in $\gtrsim 1$ s from the merger, it is also foreseen that a stable neutron star may be born 128.
@@
The ejecta emitted in connection with the merger event have different origin and composition 125; 128, as displayed in the right panel of Fig. 2. During the merger, dynamical ejecta form. These consist of shock-driven ejecta (produced during the collision of the two neutron stars, with outflows expanding across all directions and especially in the polar region) and tidal ejecta (due to mass outflows confined in the proximity of the equatorial plane and gaining sufficient angular momentum to be ejected via pre-merger tidal forces), with the tidal component being characterized by a lower electron fraction since its matter is not affected by shock heating and neutrino irradiation. More massive ejecta can originate from the unbinding of the outer layers of the newly born stable or metastable neutron star caused by the build-up of magnetic pressure.
@@
The accretion disk surrounding the central compact object can also produce an outflow, whose properties depend on the nature of the compact object, and is mainly driven by viscous heating and nuclear energy release 129. If the central compact object is a black hole, then the disk is dense enough to be initially opaque for neutrinos. As the temperature decreases, neutrino cooling tends to become inefficient, and viscous heating in the accretion disk leads to the isotropic ejection of material. If the central compact object is a neutron star, then a neutrino-driven wind expands in the proximity of the polar region, in addition to the viscous-driven disk ejecta.
@@
The observation of the gamma-ray burst GRB 170817A in connection with the gravitational wave event GW170817 50 has provided the first observational proof that the spinning black hole at the center of the merger event can harbor a Poynting-flux dominated outflow observed as short gamma-ray burst 130. Long and short gamma-ray bursts have been assumed to originate from collapsing massive stars and neutron-star mergers, respectively. However, recent observations challenge this paradigm 14; 131, with novel ideas being proposed to explain the association of kilonovae to long gamma-ray bursts 132; 133.
@@ 2.2.2 Impact of neutrino physics on multi-messenger observables
Neutrinos are fundamental in the very hot and dense regions formed following the collision of two neutron stars 8, with neutrino physics affecting the thermodynamics of the remnant as well as the composition of the outflows. The neutrino density in the core of neutron-star merger remnants is roughly comparable to the one characterizing the supernova core, although rapidly dropping in time and with an overall abundance of electron antineutrinos over neutrinos (due to the neutron richness of the environment). We expect neutrinos and antineutrinos, mostly of electron flavors, to be copiously emitted up to tens of seconds from the coalescence of two neutron stars, with an average energy of $\mathcal{O}(10)$ MeV (cf. Table 1).
@@
The detection chances of thermal neutrinos from neutron-star merger remnants are poor 134; in fact, assuming that the ratio between neutron-star mergers and core-collapse supernovae in our Galaxy is comparable to the one obtained from the local rates of such sources–see Table 1–we expect a neutron-star merger to be $10^{-4}$$10^{-2}$ less frequent than a supernova. This is also visible from the left panel of Fig. 1 where the orange lines, representing the fluence from a Galactic neutron-star merger remnant (note that the distance of $10$ kpc has been chosen for representative purposes), is contrasted with the one from a core-collapse supernova at the same distance from Earth.
@@
The modeling of neutrino flavor conversion in the core of neutron-star merger remnants is as challenging as for core-collapse supernovae, with additional complications linked to the disk geometry 80. Similar to the supernova case, fast and collisional instabilities can take place in the proximity of the neutrino decoupling region, as shown in the right panel of Fig. 2, triggering flavor conversion. In addition, because of the abundance of electron antineutrinos over neutrinos, the neutrino-neutrino self-interaction term can have opposite sign with respect to the matter term in the Hamiltonian; the cancellation between these two potentials can lead to a flavor conversion phenomenon known as matter-neutrino resonance at larger distances from the central compact object 135. However, recent work highlights the importance of the modeling of the neutrino angular distributions to reliably assess whether any flavor conversion should be expected from the matter-neutrino resonance 136.
@@
While general relativistic neutrino magnetohydrodynamic simulations of neutron-star merger remnants do not take into account neutrino flavor conversion, the latter likely has implications on the synthesis of the elements heavier than iron via the rapid neutron capture process ($r$-process). Hence, neutrino flavor conversion can affect the electromagnetic kilonova properties, modifying the production rate of lanthanides especially in the proximity of the polar region 11; 12; 137; 13.
@@
The multi-messenger data from the GW 170817 event can already be employed to place constraints on the existence of New Physics, also linked to the neutrino sector. However, as it is the case for core-collapse supernovae, it is crucial to first gain control on the standard physics involved in neutron-star mergers, especially for what concerns neutrino transport and the physics linked to the magnetohydrodynamic instabilities, which still involves several approximations 123; 138; 139.
@@ 2.2.3 Diffuse neutrino background from neutron-star mergers
To simulate the binary merger population diversity, we should rely on a large suite of long-term general-relativistic neutrino-radiation magnetohydrodynamics merger simulations. These are not yet available. Nevertheless, taking into account the very large uncertainties on the cosmic rate of neutron-star mergers 140 (which predict a rate of neutron-star mergers that is one-to-two orders of magnitude smaller than the supernova one, or even smaller, cf. Table 1) and on the neutrino fluence from each merger event (which is expected to be comparable to or smaller than the core-collapse supernova one, see the left panel of Fig. 1), the diffuse neutrino flux from these sources should be at best one-to-three orders of magnitude smaller than the DSNB one (cf. the right panel of Fig. 1), as already suggested by preliminary estimations 141. Hence, thermal neutrinos from neutron-star merger remnants cannot be adopted to probe the merger remnant population.
@@
  • $\ast$ The detection of neutrinos from a Galactic supernova will be essential to alert observers focusing on the electromagnetic spectrum about the upcoming collapse, provide information on the pre-explosion physics (complementing the input coming from gravitational waves), the nature and properties of the central compact object.

### v2

@@ 2.1.3 Detection strategies for multi-messenger searches
Among existing neutrino telescopes, the ones allowing for the largest supernova neutrino event rate are Super-Kamiokande 94 and the IceCube Neutrino Observatory 95 (cf. Table 1 of 96 for an overview on neutrino telescopes and expected event statistics). While for core-collapse supernovae in our Galaxy, IceCube would guarantee the largest statistics, Super-Kamiokande would be background free with good approximation, given the small time window characterizing the supernova neutrino burst and should allow us to reconstruct event-by-event energy information.
Within the multi-messenger framework, the SuperNova Early Warning System (SNEWS) is an operating global network of neutrino telescopes that aims to release a prompt alert of core-collapse events to enable multi-messenger observations 97. In addition to being part of SNEWS, both Super-Kamiokande and IceCube are heavily involved in the development of multi-messenger alert systems. For example, Super-Kamiokande, in collaboration with the KamLAND detector, has established an alert system that searches for pre-supernova neutrinos and could release a warning about an upcoming electromagnetic burst about a dozen hours earlier 98. The IceCube supernova data acquisition system is being updated to swiftly react to alerts from the LIGO-Virgo-KAGRA gravitational wave network 99.
Hyper-Kamiokande 100 is expected to begin taking data in 2027 and the liquid scintillator Jiangmen Underground Neutrino Observatory (JUNO) 101 should be operative already in 2025. These neutrino telescopes will guarantee large event statistics of supernova neutrinos with a few- to sub-degree pointing capabilities. If able to detect supernova neutrinos, the liquid-Argon Deep Underground Neutrino Experiment (DUNE) 102 could provide key complementarity with respect to Hyper-Kamiokande and JUNO, being mostly sensitive to electron neutrinos rather than electron antineutrinos. In addition, experiments for the direct detection of dark matter, exploiting the coherent elastic scattering of neutrinos on nuclei, such as XENONnT 103 at the Gran Sasso National Laboratory and the planned neutrino detector RES-NOVA 104, have the potential to complement large-scale neutrino detectors, being sensitive to all neutrino flavors.
@@ 2.1.4 Supernova physics signatures imprinted in the neutrino signal
Despite the dependence of the neutrino signal on the progenitor mass of the collapsing star as well as the nuclear equation of state, neutrino detection will be instrumental in exploring the following 39; 87.
  • Alert of the stellar collapse. Neutrinos reach Earth up to one day earlier than electromagnetic radiation. The early detection of the supernova neutrino burst could inform astronomers of the upcoming electromagnetic burst, enabling the electromagnetic detection of the shock breakout 105. We can also use the neutrino signal to localize the stellar collapse in the sky and define the time window to search for gravitational waves as well as build a matching filter for gravitational wave searches 106; 107. Additionally, pre-supernova neutrinos 108 produced in the late burning stages of the stellar collapse (cf. 109 for a summary of the thermal processes leading to neutrino production during the stages of stellar evolution of a massive star) could potentially provide an even earlier indication of the upcoming stellar explosion.
  • Black hole forming collapses. The temporal evolution of the neutrino signal carries distinctive signatures, ending abruptly at the time of black hole formation 110. Such collapses would not be detectable through electromagnetic radiation. Therefore, the detection of neutrinos and gravitational waves could be essential to assess the rate of black hole forming collapses, complementing ongoing electromagnetic searches 111.
  • Pre-explosion physics. The temporal evolution of the neutrino event rate (during the accretion phase) carries clear imprints of the hydrodynamical instabilities characterizing the accretion phase, such as SASI and convection 112. Rotation can also affect the neutrino signal 113; 114.
  • Neutron star properties. The late-time neutrino emission (during the Kelvin-Helmoltz cooling phase) strongly depends on the nuclear equation of state and the neutron-star properties. Despite the presence of degeneracies in the neutrino signal, its detection could place constraints on the physics of the compact remnant 115; 116.
  • Physics beyond the Standard Model. The presence of exotic signatures in the time evolution of the neutrino signal and/or its energy distribution could be a manifestation of New Physics 117. However, degeneracies among multiple New Physics scenarios exist and are yet to be fully understood 118. According to the type of exotic features eventually appearing in the observed neutrino signal, before to draw any conclusions on the discovery of New Physics relying on core-collapse supernovae, one would need to model the standard core-collapse supernova scenario (including flavor conversion) consistently–an ambitious goal yet to be achieved.
@@ 2.1.5 Diffuse supernova neutrino background
Every second a supernova explodes somewhere in the Universe. The cumulative flux of neutrinos from all core-collapse supernovae is named diffuse supernova neutrino background (DSNB) 119. This stationary and isotropic flux depends on the core-collapse supernova rate as a function of redshift, $R(z)$, and the time-integrated (anti)neutrino spectral energy distribution $F_{\nu}(E)$ emitted from each collapsing star:
@@
To date, the largest uncertainty plaguing the forecast of the DSNB signal is the core-collapse supernova rate 120, while the DSNB negligibly depends on the uncertainties on the initial mass function 121. The unknown rate of collapses leading to black hole formation 122 and magnetorotational collapses 123, as well as the metallicity evolution of the host galaxies 124, and the impact of mass transfer during binary evolution (if the collapsing star happens to be in a binary system) 125 are some of the uncertainties in the forecast of the DSNB. In addition, the unknowns linked to the supernova microphysics (such as the neutron star mass, high-density nuclear equation of state, and the flavor conversion physics 126; 127) also affect the DSNB.
The upcoming DSNB detection will be instrumental in placing neutrino-based constraints on the properties of the supernova population, the average redshift distribution of supernovae in the local Universe, and the average supernova energetics. In particular, the rate of core-collapse supernovae measured through electromagnetic surveys could be biased because electromagnetically dark collapses could be missed, but such collapses would contribute to the DNSB 128.
The Gadolinium loading of the water Cherenkov detector Super-Kamiokande has drastically enhanced the signal over background ratio, already excluding some of the most optimistic DSNB models as well as hinting towards a signal excess at $2.3\sigma$ 129; 71. The latest Super-Kamiokande upper limits on $\bar{\nu}_{e}$’s of astrophysical origin are presented in the right panel of Fig. 1 together with the expected DSNB signal. In the future, Hyper-Kamiokande 100, possibly loaded with Gadolinium, and JUNO 130 will allow us to investigate the DSNB physics with larger statistics (cf. Table 2 of 123 for estimations of the expected neutrino number of events). On the other hand, neutrino and direct-detection dark matter experiments, exploiting the coherent elastic neutrino scattering on nuclei, can place upper limits on the DSNB emission in the non-electron flavors 131.
@@ 2.2.1 Neutron-star merger physics
In 2017, the joint detection of gravitational waves and electromagnetic radiation from the merger of two neutron stars confirmed the relevance of these objects for astrophysics and fundamental physics 6; 132; see Table 1 for a summary of the multi-messenger energetics. The detection in 2017 has provided a strong indication that at least a fraction of the elements heavier than iron can be synthesized in these sources. Yet, several aspects of the physics of merger remnants remain to be clarified, such as the interplay between magnetic fields and neutrino transport, and their degree of sophistication 133; 134.
@@
Following the merger event, the accretion disk undergoes cooling via neutrino emission, with neutrino-matter interactions determining the composition of the merger outflows 135. According to the mass of the system and the nuclear equation of state, the central compact object (a newly-born neutron star) can immediately collapse into a black hole or it can be a rapidly spinning neutron star remnant. In the latter scenario, a hyper-massive 136 or supra-massive 137 neutron star may form, which is expected to collapse into a black hole in $\gtrsim 1$ s from the merger, it is also foreseen that a stable neutron star may be born 138.
@@
The ejecta emitted in connection with the merger event have different origin and composition 135; 138, as displayed in the right panel of Fig. 2. During the merger, dynamical ejecta form. These consist of shock-driven ejecta (produced during the collision of the two neutron stars, with outflows expanding across all directions and especially in the polar region) and tidal ejecta (due to mass outflows confined in the proximity of the equatorial plane and gaining sufficient angular momentum to be ejected via pre-merger tidal forces), with the tidal component being characterized by a lower electron fraction since its matter is not affected by shock heating and neutrino irradiation. More massive ejecta can originate from the unbinding of the outer layers of the newly born stable or metastable neutron star caused by the build-up of magnetic pressure.
@@
The accretion disk surrounding the central compact object can also produce an outflow, whose properties depend on the nature of the compact object, and is mainly driven by viscous heating and nuclear energy release 139. If the central compact object is a black hole, then the disk is dense enough to be initially opaque for neutrinos. As the temperature decreases, neutrino cooling tends to become inefficient, and viscous heating in the accretion disk leads to the isotropic ejection of material. If the central compact object is a neutron star, then a neutrino-driven wind expands in the proximity of the polar region, in addition to the viscous-driven disk ejecta.
@@
The observation of the gamma-ray burst GRB 170817A in connection with the gravitational wave event GW170817 57 has provided the first observational proof that the spinning black hole at the center of the merger event can harbor a Poynting-flux dominated outflow observed as short gamma-ray burst 140. Long and short gamma-ray bursts have been assumed to originate from collapsing massive stars and neutron-star mergers, respectively. However, recent observations challenge this paradigm 16; 141, with novel ideas being proposed to explain the association of kilonovae to long gamma-ray bursts 142; 143.
@@ 2.2.2 Impact of neutrino physics on multi-messenger observables
Neutrinos are fundamental in the very hot and dense regions formed following the collision of two neutron stars 10, with neutrino physics affecting the thermodynamics of the remnant as well as the composition of the outflows. The neutrino density in the core of neutron-star merger remnants is roughly comparable to the one characterizing the supernova core, although rapidly dropping in time and with an overall abundance of electron antineutrinos over neutrinos (due to the neutron richness of the environment). We expect neutrinos and antineutrinos, mostly of electron flavors, to be copiously emitted up to tens of seconds from the coalescence of two neutron stars, with an average energy of $\mathcal{O}(10)$ MeV (cf. Table 1).
@@
The detection chances of thermal neutrinos from neutron-star merger remnants are poor 144; in fact, assuming that the ratio between neutron-star mergers and core-collapse supernovae in our Galaxy is comparable to the one obtained from the local rates of such sources–see Table 1–we expect a neutron-star merger to be $10^{-4}$$10^{-2}$ less frequent than a supernova. This is also visible from the left panel of Fig. 1 where the orange lines, representing the fluence from a Galactic neutron-star merger remnant (note that the distance of $10$ kpc has been chosen for representative purposes), is contrasted with the one from a core-collapse supernova at the same distance from Earth.
@@
The modeling of neutrino flavor conversion in the core of neutron-star merger remnants is as challenging as for core-collapse supernovae, with additional complications linked to the disk geometry 89. Similar to the supernova case, fast and collisional instabilities can take place in the proximity of the neutrino decoupling region, as shown in the right panel of Fig. 2, triggering flavor conversion. In addition, because of the abundance of electron antineutrinos over neutrinos, the neutrino-neutrino self-interaction term can have opposite sign with respect to the matter term in the Hamiltonian; the cancellation between these two potentials can lead to a flavor conversion phenomenon known as matter-neutrino resonance at larger distances from the central compact object 145. However, recent work highlights the importance of the modeling of the neutrino angular distributions to reliably assess whether any flavor conversion should be expected from the matter-neutrino resonance 146.
@@
While general relativistic neutrino magnetohydrodynamic simulations of neutron-star merger remnants do not take into account neutrino flavor conversion, the latter likely has implications on the synthesis of the elements heavier than iron via the rapid neutron capture process ($r$-process). Hence, neutrino flavor conversion can affect the electromagnetic kilonova properties, modifying the production rate of lanthanides especially in the proximity of the polar region 13; 14; 147; 15.
@@
The multi-messenger data from the GW 170817 event can already be employed to place constraints on the existence of New Physics, also linked to the neutrino sector. However, as it is the case for core-collapse supernovae, it is crucial to first gain control on the standard physics involved in neutron-star mergers, especially for what concerns neutrino transport and the physics linked to the magnetohydrodynamic instabilities, which still involves several approximations 133; 148; 149.
@@ 2.2.3 Diffuse neutrino background from neutron-star mergers
To simulate the binary merger population diversity, we should rely on a large suite of long-term general-relativistic neutrino-radiation magnetohydrodynamics merger simulations. These are not yet available. Nevertheless, taking into account the very large uncertainties on the cosmic rate of neutron-star mergers 150 (which predict a rate of neutron-star mergers that is one-to-two orders of magnitude smaller than the supernova one, or even smaller, cf. Table 1) and on the neutrino fluence from each merger event (which is expected to be comparable to or smaller than the core-collapse supernova one, see the left panel of Fig. 1), the diffuse neutrino flux from these sources should be at best one-to-three orders of magnitude smaller than the DSNB one (cf. the right panel of Fig. 1), as already suggested by preliminary estimations 151. Hence, thermal neutrinos from neutron-star merger remnants cannot be adopted to probe the merger remnant population.
@@
  • $\ast$ The detection of neutrinos from a Galactic supernova will be essential to alert observers focusing on the electromagnetic spectrum about the upcoming collapse, provide information on the pre-explosion physics (complementing the input coming from gravitational waves), as well as the nature and properties of the central compact object.